# explExample

### v2

@@ 4 LRNR approximation of classical solutions
In this section, we will discuss how a LRNR can be used to approximate classical solutions to (2.1). In this case, the solution is given by the method of characteristics, and is known to have an efficient approximation using transported subspaces [71]. For the classical case, it is simpler to suspend the assumption that the initial condition is piecewise constant with finitely many jumps (that is, $u_{0}\in\overline{\mathcal{U}}$). Then, for a smooth initial condition $u_{0}$ the classical solution is
@@
Due to the absence of rarefaction and shock waves in the classical case, the solutions simplifies in two major ways: (1) The inverse $(\cdot)^{-1}$ is used in the solution (4.1) rather than the left-inverse $(\cdot)^{+}$ used in (2.25, 2.33), and (2) the initial condition $u_{0}$ and the characteristic curves $X$ are directly used instead of their extensions (i.e. $\hat{u}_{0}$, $\widehat{X}$ or $\widecheck{X}$).
@@
[…]
@@
We will recall the transported subspaces, and discuss how the inverse that appears in the composition can be avoided straightforwardly using a trick we call the inverse-bias trick. Starting with a given transported subspace in its original form, we will first show that one can rewrite it in a certain discrete form using this trick. Next, we will show that a transported subspace in this discrete form is equivalent to a 2-layer LRNR of comparable complexity. Finally, we will approximate the classical solution to the scalar conservation law using a 2-layer LRNR. The techniques developed in this section for approximating the classical solutions will be used throughout the subsequent Section 5 that addresses the entropy solution. We refer the reader to Appendix A for a summary of notions including the transported subspaces, the inverse-bias trick, and the LRNR approximation of the classical solution introduced in this section.
@@ 4.1 The transported subspaces and the inverse-bias trick
We will review the definition of transported subspaces and introduce an approximation technique which allows one to avoid calculating the left-inverses that typically appear in method of characteristics, e.g. in (2.34). Then we will derive a useful estimate for transport subspaces when this technique is applied.
@@
Given a domain $\Omega_{x}\subset\mathbb{R}$ and a set of $r$ functions $\Phi_{r}=(\phi_{i})_{i=1}^{r}\subset P_{1}(\mathbb{R})\cap C(\mathbb{R})$ we say a set of coefficients $A\subset\mathbb{R}^{r}$ is admissible for $\Phi_{r}$ on $\Omega_{x}$ if
@@
and two sets of coefficients, $A\subset\mathbb{R}^{r}$ admissible for $\Phi_{r}$ on $\Omega_{x}$ and $B\subset\mathbb{R}^{r}$ bounded.
@@
(4.5) $\bar{H}_{r}$ $:=(B\cdot\Psi_{r})\circ(A\cdot\Phi_{r})^{-1}$
=$=\left\{\bar{h}\in P_{0}(\Omega_{x})\cap BV(\Omega_{x})\ |\ \bar{h}=f\circ g^{-1},\ f\in B\cdot\Psi_{r},\ g\in A\cdot\Phi_{r}\right\}.$
@@
(4.8) $\rho(g^{-1}(x)-x_{k})$ =$=\begin{cases}1&\text{ if }g^{-1}(x)\geq x_{k},\\ 0&\text{ if }g^{-1}(x)<x_{k},\end{cases}$
=$=\begin{cases}1&\text{ if }x\geq g(x_{k}),\\ 0&\text{ if }x<g(x_{k}),\end{cases}$
@@
This trick allows us to avoid maintaining the invertibility of $g$, as enforced by the admissibility conditions in Definition 4.1. Note how this has been a source of difficulties discussed in similar approaches [38, 9, 78]. In contrast, the functional form of LRNR is well-defined even without enforcing the invertibility of the transport map $g$.
Although the inverse-bias trick might appear to apply solely in the single dimensional setting, it extends to multiple dimensions when one views the neural network approximation as a superposition of ridgelets, for instance via the Radon transform [57, 35, 7, 68, 69]. Close connections have been made to neural networks, e.g. see [11, 61, 81].
The following lemma illustrates that the inverse-bias trick can be extended to the setting where the approximating function is continuous piecewise linear, that is, where the jump function $\rho$ is replaced by a continuous piecewise linear approximation $\rho_{\varepsilon}$. Throughout, we will make use of the piecewise linear approximation $\rho_{\varepsilon}$ of the threshold function $\rho$,
@@
Then we have that, for parameter $\varepsilon>0$ sufficiently small depending on the grid width,
@@ 4.2 LRNR approximation of transported subspaces
In this section, we will show that a 2-layer version of LRNR, a simple special case of the general definition of LRNRs introduced in (3), is able to approximate any transported subspace. The rank of the approximating LRNR is equivalent to the dimension of the transported subspace.
{expl}
[A 2-layer version of LRNR] Suppose we are given
@@
a pair of coefficient sets $(A,B)$, and a given parameter $\varepsilon>0$, there exists a LRNR $H_{r}$ that has d.o.f $K\lesssim\sum_{i=1}^{r}N_{i}$, width $M\sim K$, and depth $L=2$, along with coefficient sets $C,D$ and an affine map $\mu:A\times B\to C\times D$ mapping $\mu(\alpha,\beta)=(\gamma,\theta),$ such that corresponding members $\bar{h}(\alpha,\beta)\in\bar{H}_{r-1}$ and $h(\gamma,\theta)\in H_{r}$ satisfy
@@ 4.3 LRNR approximation of classical solutions
Now that we have shown that there are LRNR approximations of transported subspaces (Theorem 4.5), it is natural to ask whether the classical solution can be approximated uniformly well by a LRNR. We show that this is indeed the case.
@@
there is a LRNR $H_{r}$ with rank $r=3$, d.o.f $K$, width $M\sim K$, and depth $L=2$ for every given $\varepsilon>0$ such that, for coefficients $\gamma(t),\theta(t)$ that are linear or constant functions of $t$, the member $h(t)=h(\gamma(t),\theta(t))\in H_{r}$ achieves the uniform error
@@
We remark here that the classical norm $\lVert{u_{0}^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}$ can be avoided, as will become apparent in the results in the next section (Theorem 5.7). Its presence here is attributable to the separate approximation of the initial condition $u_{0}$ and the characteristic curves $X$, without taking into consideration the simplifications arising in their composition. Still, this result uses a straightforward approximation and demonstrates that, given sufficient width, the intrinsic low dimension can be exploited by LRNR.
We briefly emphasize again that while the transported subspace requires that the coefficients be admissible, whereas LRNR does not require similar constraints. The representation makes sense for any coefficients $\gamma$ and $\theta$, and this is a consequence of the inverse-bias trick.
@@ 5 LRNR approximation of entropy solutions
Having constructed an efficient LRNR approximation of the classical solution, we turn our attention to the LRNR approximation of the entropy solution. As briefly discussed in the beginning of Section 4, there are important differences between the classical solution and the entropy solution. Let us explain how these differences introduce significant challenges in constructing the LRNR approximation.
Recall from Section 2 that we have introduced two new representations of the entropy solution, namely, one using the rarefied characteristics $\widehat{X}$ (2.25) and another using the relief characteristics $\widecheck{X}$ (2.33):
@@
The rarefied characteristics $\widehat{X}$ is a direct generalization of the classical chracteristics that allows representing shocks and rarefaction waves. It is monotone, thus the inverse-bias trick can still be applied directly to handle the left-inverse $(\cdot)^{+}$. However, as we discuss below, the rarefied characteristics are not in a compositional low rank form that is compatible with an efficient representation by LRNRs. On the other hand, the relief chracteristics $\widecheck{X}$ do have a compositional low rank structure (as discussed in Section 3) that is compatible with LRNRs. However, it is not monotone and hence the left-inverse cannot be removed using the inverse-bias trick directly.
Therefore, our goal here is to devise a LRNR approximation that agrees with the rarefied characteristics $\widehat{X}$, while simultaneously exploiting the compositional low-rank structure of the relief characteristics $\widecheck{X}$. We build a construction by decomposing the relief characteristics into sum of local-in-time contributions, and introduce a LRNR structure per each summand via the same technique used to for relief characteristics. This allows us to take advantage of the compositional structure while approximately maintaining monotonicity.
In this section we will address these challenges involved and construct an efficient LRNR approximation of the entropy solution. We first provide a broad outline in Section 5.1, then introduce various discretizations that gradually reveal the neural-network-like structure in Sections 5.2 and 5.3, to finally work towards the main Theorem 5.7 in Section 5.4. Throughout, we refer the reader to Appendix A for a glossary of various approximations introduced in this section.
@@ 5.1 Outline of approach
Figure 7: The time evolution of the rarefied characteristics $\widehat{X}(\cdot,t)$ for the stationary shock example in Figure 5, zoomed into the region relevant to shock propagation. The constant interval (marked by $\Upsilon_{t}$) expands over time $0<t_{1}<t_{2}<t_{3}$. The two kinks at the endpoints (marked by two arrows) of this interval travel in opposite directions over time.
In this section, we will motivate and outline our general approximation approach. We begin by detailing the technical challenges mentioned in the introduction of this section. In light of the inverse-bias trick in Section 4.1, it might appear that the approximation of the entropy solution is now a simple matter of approximating the rarefied characteristics $\widehat{X}$ (2.22). Unfortunately, the rarefied characteristics manifold given by
@@
itself has a slowly decaying Kolmogorov $n$-width in the presence of shocks. For example, recall the stationary shock example from Figure 5. Observe how the rarefied characteristics $\widehat{X}(\cdot,t)$ evolves over time, as illustrated in Figure 7. The constant region of $\widehat{X}(\cdot,t)$ describing the shock expand with time, according to the Rankine-Hugoniot condition, and this leads to kinks that travel across the domain $\widehat{\Omega}_{x}$. As a result, the $n$-width in $L^{2}$-norm must be $\gtrsim n^{-3/2}$ [83, 60, 31, 72]. This is an obstruction to the strategy of finding a compositional reduced model directly, by using e.g. transported subspaces (Definition 4.2). In other words, there is now a slow decay in the $n$-width of both the solution manifold and the rarefied characteristics manifold. [^2] To provide a succinct analogy: The classical characteristics are to the solution manifold, what the shock time function is to the rarefied characteristics manifold.
One natural question to ask is whether it is possible to construct compositional reduced models yet again, this time for the manifold of rarefied characteristics $\mathcal{M}_{\widehat{X}}$. A low-rank structure is apparent in the relief characteristics $\widecheck{X}$ (2.32), in the superpositions of a few basis functions, as discussed in Section 3; see the diagram in (3.1). Still, we need care in introducing the shock-related jump term $\rho(\lambda(x)-t)$ that appears in the inner-most composition (2.32). In the relief characteristics, this term is used in conjunction with the left-inverse to replace the portions of the rarefied characteristic curves with constant regions. LRNRs are defined using ReLU activations so they are continuous piecewise linear functions; as such, the jump discontinuity must be approximated carefully so that the low-rank compositional structure is preserved. We propose a LRNR approximation that satisfies these requirements, outlined as follows.
@@
    (5.5) $\widehat{X}(\cdot,t)\approx\barwedge(\cdot,t)\approx\hat{I}(\cdot)+tv(\cdot)+\sum_{\mathclap{\begin{subarray}{c}\text{partial sum}\\ \text{over }k\end{subarray}}}t\,\eta_{k}\left(\,\cdot\,\right),\qquad\lVert{t\eta_{k}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\lesssim\varepsilon.$
@@
We present our construction as follows. In Section 5.2, we derive a time-layered approximation of rarefied characteristics $\widehat{X}$ that is valid for short time intervals, then in Section 5.3 we construct a discretization of the form (5.6), then finally in Section 5.4 we present the LRNR approximation result.

### v4

@@ 4 LRNR approximation of classical solutions
In this section, we will discuss how a LRNR can be used to approximate classical solutions to (2.1). In this case, the solution is given by the method of characteristics (2.30), and is known to have an efficient approximation using transported subspaces [71]. For the classical case, it is mandatory to suspend the assumption that the initial condition is piecewise constant with finitely many jumps (that is, $u_{0}\in\overline{\mathcal{U}}$). Then, for a smooth initial condition $u_{0}\in\mathcal{U}$ (2.2) the classical solution is
@@
Due to the absence of rarefaction and shock waves in the classical case, the solutions simplifies in two major ways: (1) The inverse $(\cdot)^{-1}$ is used in the solution (4.1) rather than the left-inverse $(\cdot)^{+}$ used in (2.25, 2.32), and (2) the initial condition $u_{0}$ and the characteristic curves $X$ are directly used instead of their extensions (i.e. $\hat{u}_{0}$, $\widehat{X}$ or $\widecheck{X}$).
@@
 $\qquad\text{ with }r=3\quad\text{ and }\quad\left\{\begin{aligned} \gamma^{1}_{1}&=1,&\gamma^{1}_{2}&=t,&\gamma^{1}_{3}&=0,\\ \varphi_{1}&=\operatorname{Id}&\varphi_{2}&=F^{\prime}\circ u_{0},&\\ \gamma^{2}_{1}&=1,&\gamma^{2}_{2}&=0,&\gamma^{2}_{3}&=0,\\ \phi_{1}&=u_{0}.\end{aligned}\right$.
@@
We will recall the transported subspaces, and discuss how the inverse that appears in the composition can be avoided straightforwardly using a trick we call the inverse-bias trick. Starting with a given transported subspace in its original form, we will first show that one can rewrite it in a certain discrete form using this trick. Next, we will show that a transported subspace in this discrete form is equivalent to a 2-layer LRNR of comparable complexity. Finally, we will approximate the classical solution to the scalar conservation law using a 2-layer LRNR. The techniques developed in this section for approximating the classical solutions will be used throughout the subsequent Sec. 5 that addresses the entropy solution. We refer the reader to Appendix A for a summary of notions including the transported subspaces, the inverse-bias trick, and the LRNR approximation of the classical solution introduced in this section.
@@ 4.1 The transported subspaces and the inverse-bias trick
We will review the definition of transported subspaces and introduce an approximation technique which allows one to avoid calculating the left-inverses that typically appear in method of characteristics, e.g. in (2.30, 2.33). Then we will derive a useful estimate for transport subspaces when this technique is applied.
@@
Given a domain $\Omega_{x}\subset\mathbb{R}$ and a set of $r$ functions $\Phi_{r}=(\phi_{i})_{i=1}^{r}\subset P_{1}(\mathbb{R})\cap C(\mathbb{R})$ we say a coefficient set $A\subset\mathbb{R}^{r}$ is admissible for $\Phi_{r}$ on $\Omega_{x}$ if
@@
and two coefficient sets, $A\subset\mathbb{R}^{r}$ admissible for $\Phi_{r}$ on $\Omega_{x}$ and $B\subset\mathbb{R}^{r}$ bounded.
@@
(4.5) $\bar{H}_{r}$ $=(B\cdot\Psi_{r})\circ(A\cdot\Phi_{r})^{-1}$
$=\left\{\bar{h}\in P_{0}(\Omega_{x})\cap BV(\Omega_{x})\ |\ \bar{h}=f\circ g^{-1},\ f\in B\cdot\Psi_{r},\ g\in A\cdot\Phi_{r}\right\}.$
@@
[…]
@@
Careful consideration in Lem. 4.3 below shows that the inverse bias trick applies to non-decreasing functions that have a left-inverse but not necessarily an inverse. While in this paper analytical expression for $g$ originate from hyperbolic conservation laws and are thus left-invertible, we do not require invertibility of their numerical approximations $g_{\varepsilon}$. Indeed, since the right hand side of the inverse bias trick is independent of the inverse, we have $\rho(g^{-1}(x)-x_{k})=\rho(x-g(x_{k})\approx\rho(x-g_{\varepsilon}(x))$ for which we can easily control the error without assuming invertibility of $g_{\varepsilon}$, as e.g. in Lem. 4.3. In principle, one can further apply the right hand side to arbitrary non-invertible functions $g$ such as multi-valued characteristics at shocks, but this would invalidate the left hand side and with it the original meaning of the identity. For such cases, we require additional techinques, as we discuss in Sec. 5. Note that this issue of invertibility has been a source of difficulties discussed in related approaches [38, 9, 78].
In the following lemma the approximating function is continuous and piecewise linear; therein the jump function $\rho$ is replaced by a continuous piecewise linear approximation $\rho_{\varepsilon}$ defined as
@@
Then we have for parameter $\varepsilon>0$ sufficiently small depending on the grid width,
@@
Although the inverse-bias trick might appear to apply solely in the single dimensional setting, it extends to multiple dimensions when one views the neural network approximation as a superposition of ridge functions, for instance via the Radon transform [57, 35, 7, 68, 69]. Close connections have been made to neural networks, e.g. see [11, 61, 81].
@@ 4.2 LRNR approximation of transported subspaces
In this section, we will show that a 2-layer version of LRNR, a simple special case of the general definition of LRNRs introduced in Sec. 3, is able to approximate any transported subspace. The rank of the approximating LRNR is equivalent to the dimension of the transported subspace.
Example 4.5 (A 2-layer version of LRNR).
Suppose we are given
@@
a pair of coefficient sets $(A,B)$, and a given parameter $\varepsilon>0$, there exists a LRNR $H_{r}$ that has d.o.f $K\lesssim\sum_{i=1}^{r}N_{i}$, width $M\sim K$, and depth $L=2$, along with coefficient sets $C,D$ and an affine map $\mu:A\times B\to C\times D$ mapping $\mu(\alpha,\beta)=(\gamma,\theta),$ such that the members $\bar{h}(\alpha,\beta)\in\bar{H}_{r-1}$ and $h(\gamma,\theta)\in H_{r}$ satisfy
@@ 4.3 LRNR approximation of classical solutions
Now that we have shown that there are LRNR approximations of transported subspaces (Thm. 4.6), it is natural to ask whether the classical solution can be approximated uniformly well by a LRNR. We show that this is indeed the case.
@@
there is a LRNR $H_{r}$ with rank $r=3$, d.o.f $K$, width $M\sim K$, and depth $L=2$ such that, for coefficients $\gamma(t),\theta(t)$ that are linear or constant functions of $t$, the member $h(t)=h(\gamma(t),\theta(t))\in H_{r}$ achieves the uniform error
@@
We remark here that the classical norm $\lVert{u_{0}^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}$ can be avoided, as will become apparent in the results in the next section (Thm. 5.7). Its presence here is attributable to the separate approximation of the initial condition $u_{0}$ and the characteristic curves $X$, without taking into consideration the simplifications arising in their composition. Still, this result uses a straightforward approximation and demonstrates that, given sufficient width, the intrinsic low dimension can be exploited by LRNR.
We briefly emphasize again that while the transported subspace requires that its coefficient sets be admissible, the LRNR does not require similar constraints. The representation makes sense for any coefficients $\gamma$ and $\theta$, and this is a consequence of the inverse-bias trick.
@@ 5 LRNR approximation of entropy solutions
Having constructed an efficient LRNR approximation of the classical solution, we turn our attention to the LRNR approximation of the entropy solution. As briefly discussed in the beginning of Sec. 4, there are important differences between the classical solution and the entropy solution. Let us explain how these differences introduce significant challenges in constructing the LRNR approximation.
Recall from Sec. 2 that we have introduced two new representations of the entropy solution, namely, one using the rarefied characteristics $\widehat{X}$ (2.25) and another using the relief characteristics $\widecheck{X}$ (2.32):
@@
The rarefied characteristics $\widehat{X}$ is a direct generalization of the classical chracteristics that allows representing shocks and rarefaction waves. It is monotone, thus the inverse-bias trick can still be applied directly to handle the left-inverse $(\cdot)^{+}$. However, as we discuss below, the rarefied characteristics are not in a compositional low rank form that is compatible with an efficient representation by LRNRs. On the other hand, the relief chracteristics $\widecheck{X}$ do have a compositional low rank structure (as discussed in Sec. 3) that is compatible with LRNRs. However, it is not monotone and hence the left-inverse cannot be removed using the inverse-bias trick directly and an additional layer of approximation is required (see paragraphs following (4.7)).
Therefore, our goal here is to devise a LRNR approximation that agrees with the rarefied characteristics $\widehat{X}$, while simultaneously exploiting the compositional low-rank structure of the relief characteristics $\widecheck{X}$. We build a construction by decomposing the rarefied characteristics into sum of local-in-time contributions, and introduce a LRNR structure per each summand via the same technique used to for relief characteristics. This allows us to take advantage of the compositional structure while not necessarily maintaining monotonicity.
In this section we will address these challenges involved and construct an efficient LRNR approximation of the entropy solution. We first provide a broad outline in Sec. 5.1, then introduce various discretizations that gradually reveal the neural-network-like structure in Sec. 5.2 and Sec. 5.3, to finally work towards the main Thm. 5.7 in Sec. 5.4. Throughout, we refer the reader to Appendix A for a glossary of various approximations introduced in this section.
@@ 5.1 Outline of approach
Figure 7: The time evolution of the rarefied characteristics $\widehat{X}(\cdot,t)$ for the stationary shock example in Fig. 5, zoomed into the region relevant to shock propagation. The constant interval (marked by $\Upsilon_{t}$) expands over time $0<t_{1}<t_{2}<t_{3}$. The two kinks at the endpoints (marked by two arrows) of this interval travel in opposite directions over time.
In this section, we will motivate and outline our general approximation approach. We begin by detailing the technical challenges mentioned in the introduction of this section. In light of the inverse-bias trick in Sec. 4.1, it might appear that the approximation of the entropy solution is now a simple matter of approximating the rarefied characteristics $\widehat{X}$ (2.22). Unfortunately, the rarefied characteristics manifold given by
@@
itself has a slowly decaying Kolmogorov $n$-width in the presence of shocks. For example, recall the stationary shock example from Fig. 5. Observe how the rarefied characteristics $\widehat{X}(\cdot,t)$ evolves over time, as illustrated in Fig. 7. The constant region of $\widehat{X}(\cdot,t)$ describing the shock expand with time, according to the Rankine-Hugoniot condition, and this leads to kinks that travel across the domain $\widehat{\Omega}_{x}$. As a result, the $n$-width in $L^{2}$-norm must be $\gtrsim n^{-3/2}$ [83, 60, 31, 72]. This is an obstruction to the strategy of finding a compositional reduced model directly, by using e.g. transported subspaces (Def. 4.2). In other words, there is now a slow decay in the $n$-width of both the solution manifold and the rarefied characteristics manifold. [^2] To provide a succinct analogy: The classical characteristics are to the solution manifold, what the shock time function is to the rarefied characteristics manifold.
One natural question to ask is whether it is possible to construct compositional reduced models yet again, this time for the manifold of rarefied characteristics $\mathcal{M}_{\widehat{X}}$. A low-rank structure is apparent in the relief characteristics $\widecheck{X}$ (2.31), in the superpositions of a few basis functions, as discussed in Sec. 3; see the diagram in (3.1). Still, we need care in introducing the shock-related jump term $\rho(\lambda(x)-t)$ that appears in the inner-most composition (2.31). In the relief characteristics, this term is used in conjunction with the left-inverse to replace the portions of the rarefied characteristic curves with constant regions. LRNRs are defined using ReLU activations so they are continuous piecewise linear functions; as such, the jump discontinuity must be approximated carefully so that the low-rank compositional structure is preserved. We propose a LRNR approximation that satisfies these requirements, outlined as follows.
@@
    (5.5) $\widehat{X}(\cdot,t)\approx\barwedge(\cdot,t)\approx\hat{I}(\cdot)+tv(\cdot)+\sum_{\mathclap{\begin{subarray}{c}\text{partial sum}\\ \text{over }k\end{subarray}}}t\,\eta_{k}\left(\,\cdot\,\right),\qquad\lVert{t\eta_{k}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\lesssim\varepsilon,$
@@
We present our construction as follows. In Sec. 5.2, we derive a time-layered approximation of rarefied characteristics $\widehat{X}$ that is valid for short time intervals, then in Sec. 5.3 we construct a discretization of the form (5.6), then finally in Sec. 5.4 we present the LRNR approximation result.